OpenStudy (anonymous):

Evaluate the formula A = pie r^2 for pie= 22 over 7 and r= 7 over 4 .

OpenStudy (mathstudent55):

\( A = \pi r^2\) \( \pi = \dfrac{22}{7}\); \(r = \dfrac{7}{4} \) \( A = \dfrac{22}{7} \times (\dfrac{7}{4})^2 \) \( A = \dfrac{22}{7} \times \dfrac{49}{16} \) \(A = \dfrac{22}{1} \times \dfrac{7}{16} \) \(A = \dfrac{11}{1} \times \dfrac{7}{8} \) \(A = \dfrac{77}{8} = 9.625\)